There is an opening for 5 positions in a company and some females and males apply for the job, 4 females and 6 males were selected for an on-site interview. based on these information calculate the following parts:

  • a) name the distribution (4 points)
  • b) What is the probability that all 4 females and just 1 male get the job offers? (8 points)
  • c) what is the probability that 2 females and 3 males get the job offers? (8 points)
Answer #1

a)

this is hypergoemtric distribution with parameter N=10 and sample size n=5

b )

P( 4 females and just 1 male get the job offers ) =(4C4)*(6C1)/(10C5) =1*6/252=1/42

c)

P(2 female and 3 male) =(4C2)*(6C3)/(10C5) =6*20/252 =10/21